OpenStudy (anonymous):

show that (gh)^-1 = (h^-1)(g^-1)

OpenStudy (anonymous):

True :)

OpenStudy (anonymous):

1 = 1

OpenStudy (anonymous):

Or, 1/gh= 1/hg

OpenStudy (anonymous):

yeah. but how do i show it?

OpenStudy (anonymous):

(gh)^-1 = 1/(gh) = (1/g) * (1/h) = (g^-1)(h^-1) = (h^-1)(g^-1) = RHS

OpenStudy (anonymous):

oh great thanks!

OpenStudy (anonymous):

oh heck no

OpenStudy (anonymous):

>.>

OpenStudy (anonymous):

this \[f^{-1}\] means the inverse function, not the reciprocal

OpenStudy (anonymous):

True, True, :)

OpenStudy (anonymous):

you are trying to show that \[(h\circ g)^{-1}=g^{-1}\circ h^{-1}\]

OpenStudy (anonymous):

it is always true that \[(ab)^{-1}=b^{-1}a{-1}\] now matter what a and be are. functions, anything. check as follows

OpenStudy (anonymous):

first of all what does \[(h\circ g)^{-1}\] mean? it means if you compose this with \[(h\circ g)\] you should get the identity function back

OpenStudy (anonymous):

in other words you know that \[(h\circ g)^{-1}\circ (h\circ g) = I\]

OpenStudy (anonymous):

oh but h and g are elements of a group. it's not a function.

OpenStudy (anonymous):

so now lets try to show that \[(h\circ g)^{-1}=g^{-1}\circ h^{-1}\] by checking that it is the right thing \[g^{-1}\circ h^{-1}\circ h \circ g =g^{-1}eg=g^{-1}g=e\]\]

OpenStudy (anonymous):

fine a group it is and i got rid of the circle notation and used e as the identity. same idea exactly

OpenStudy (anonymous):

point is if the group is not commutative, you cannot switch around the inverses. so \[(ab)^{-1}=b^{-1}a^{-1}\]

OpenStudy (anonymous):

easy in this notation anyway. \[b^{-1}a^{-1}ab=b^{-1}eb=b^{-1}b=e\]

OpenStudy (anonymous):

and likewise on the other side

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

by the way they could be functions right? permutations, anything. that is why i said above that this is completely general

OpenStudy (anonymous):

oh right okay. understood. thanks